# The Kobayashi-Hitchin correspondence for nef and big classes

### 11 Mar 2026 v1

@@ 1 Introduction
The Kobayashi-Hitchin correspondence asserts that a holomorphic vector bundle $E$ over a compact complex manifold $X$ is slope polystable with respect to a Kähler class $\{\omega\}$ on $X$ if and only if $E$ admits an $\omega$-Hermitian-Yang-Mills metric. If $\omega$ is a smooth Kähler metric, this correspondence was proved by Donaldson [Don85] and Uhlenbeck-Yau [UY86] for holomorphic vector bundles on compact Kähler manifolds, by Bando-Siu [BS94] for reflexive sheaves on compact Kähler manifolds by introducing the notion of admissible Hermitian-Yang-Mills metrics, and by Xuemiao Chen [Chen25] for reflexive sheaves on compact normal Kähler varieties. More recently, the correspondence has been extended to compact klt Kähler varieties endowed with orbifold Kähler metrics motivated by the study of the equality cases of the Bogomolov-Gieseker inequality and the Miyaoka-Yau inequality of klt varieties [Faulk22], [FO25] (refer to the introduction of [IJZ25] for more detail). There are also results when Kähler metrics have conical singularities [Li00], [JL25].
@@
Our approach is based on the Uhlenbeck compactness theorem for Yang-Mills connections, which is inspired by the approach of [Chen25]. We are also inspired by the arguments in [KO25].
@@ 1.1.1 Definitions
Our main result of this paper is Theorem 1.4. To formulate Theorem 1.4, we firstly introduce the notion of adapted closed positive $(1,1)$-currents, inspired by the notion of tame approximations of [CCHTST25] for closed positive $(1,1)$-currents with bounded potentials (refer to Definition 2.18 for singular settings).
Definition 1.1 (=Definition 2.15, Definition 2.18).
@@ 3. $(3)$
We will see that any nef and big class contains an adapted current (Proposition 2.16). One of the important examples of adapted currents is a singular Kähler-Einstein metric on a compact normal Kähler variety with log terminal singularities (Theorem 2.20). We next introduce the notion of adapted Hermitian-Yang-Mills metrics as admissible Hermitian-Yang-Mills metrics with additional two conditions, called as adapted conditions (refer to Definition 2.22 for the definition in singular settings):
Definition 1.2 (=Definition 2.21, Definition 2.22).
@@ 2. $(2)$
Let $X$ be a compact normal analytic variety and $\alpha\in H^{1,1}_{BC}(X)$ be a nef and big class. Let $\mathcal{E}$ be a reflexive sheaf on $X$. Then the following conditions are equivalent.
@@
    $\mathcal{E}$ admits a $T$-adapted Hermitian-Yang-Mills metric (refer to Definition 2.22).
Furthermore, if $\mathcal{E}$ admits a $T$-adapted Hermitian-Yang-Mills metric, then it is unique up to constant multiplications.
@@
By [CCHTST25], it has proved that a singular Kähler-Einstein metric on a compact klt variety is an adapted current. Hence we obtain the following result as a direct consequence of Corollary 1.6. Following Corollary 1.8 itself is a new result to the best of our knowledge.
@@
    $\mathcal{E}$ admits an $\omega$-adapted Hermitian-Yang-Mills metric.
Furthermore, if $\mathcal{E}$ admits an $\omega$-adapted Hermitian-Yang-Mills metric, then it is unique up to constant multiplications.
@@ 1.1.3 A priori estimates and applications
Corollary 1.9 (=Proposition 3.15, Corollary 3.30).
@@ 3. $(3)$
Let $X$ be a compact normal analytic variety and $\alpha\in H^{1,1}_{BC}(X)$ be a nef and big class. Then, for any $\alpha^{n-1}$-slope polystable reflexive sheaves $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$, their direct sum $\mathcal{E}_{1}\oplus\mathcal{E}_{2}$, tensor product $\mathcal{E}_{1}\otimes\mathcal{E}_{2}$ and wedge product $\mathcal{E}_{1}\wedge\mathcal{E}_{2}$ are $\alpha^{n-1}$-slope polystable.
@@ 1.1.4 Further discussions
If a nef and big class $\alpha$ is a Kähler class, then above map is bijective and it is the non-abelian Hodge correspondence in the case that a Higgs structure is trivial (refer to [ZZZ26] for more detailed and precise statement). Then it is a natural question whether the above map is bijective in this nef and big class-setting. The identification of the image of the above map will be important.
@@
$\sup_{X\setminus D}|s_{D}\Psi_{\infty}^{\pm 1}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}<\infty,$ (1.3)
where $h_{D}$ is a smooth hermitian metric on $\mathcal{O}(D)$ such that $|s_{D}|_{h_{D}}\leq 1$ (refer to Lemma 3.7). In fact, such $\Psi_{\infty}$ defines a hermitian metric $h_{0}\Psi_{\infty}^{2}$ on $E|_{X\setminus D}$ and it gives a $T$-adapted HYM metric (see also Proposition 3.17). In the construction of $\Psi_{\infty}^{\pm 1}$ satisfying (1.3), the mean value type inequality of Guo-Phong-Sturm [GPS24] (Lemma 3.6) plays an important role. In fact, we can easily shows
@@
holds with universal constant $C>0$ (Lemma 3.8). Then, we can prove that if the $L^{1}$ norm in LHS in (1.4), denoted by $l_{i}$, diverges to $\infty$ in $i\to\infty$, then the normalized morphism $u_{i}:=l_{i}^{-1}s_{i}$ converges to a nontrivial limit $u_{\infty}$ with ${\rm{rk}}(u_{\infty})<{\rm{rk}}E$ (Lemma 3.9). Then, as the argument in [Sim88, Section 5], we can see that the projections to subspaces of the kernel of $u_{\infty}$ define $T$-weak holomorphic subbundles of $E$. Then, if we use the Chern-Weil formula for $T$-weak holomorphic subbundles (Proposition 1.5), we can see that some of the projections produce a destabilizing subsheaf of $E$ (Lemma 3.9). It contradicts to the $\alpha^{n-1}$-slope stability of $E$. Hence the $L^{1}$-norm of (1.4) is uniformly bounded and then $\Psi_{i}$ and $\Psi_{i}^{-1}$ converge to nontrivial morphisms $\Psi_{\infty}$ and $\Phi_{\infty}$ respectively satisfying (1.3) (Lemma 3.7 and Lemma 3.8). Then we can see that $\Phi_{\infty}=\Psi_{\infty}^{-1}$ and $\Psi_{\infty}$ gives a $T$-adapted HYM metric $h_{0}\Psi_{\infty}^{2}$ on $E$ (Theorem 3.17).
@@
for some $l\in\mathbb{Z}_{>0}$ where $h_{D}$ is a smooth hermitian metric on $\mathcal{O}(D)$ with $|s_{D}|_{h_{D}}\leq 1$. The first technical lemma is to show that $g$ is holomorphic, that is, the following holds (Claim 3.27)
@@
on $X\setminus D$. [2nd step] Then, by Proposition 1.10, we obtain that a $T$-adapted HYM metric on an $\alpha^{n-1}$-slope stable vector bundle is unique (Lemma 3.26). [3rd step] Then, by applying the existence of $T$-adapted HYM metrics on $\alpha^{n-1}$-slope stable sheaves and 2nd step, we can prove that a holomorphic vector bundle $E$ admitting a $T$-adapted HYM metric is $\alpha^{n-1}$-slope polystable (Theorem 3.28). [4th step] Let $E|_{X\setminus D}=(\bigoplus_{i=1}^{k}\mathcal{E}_{i})|_{X\setminus D}$ be the decomposition into $\alpha^{n-1}$-slope stable torsion free sheaves of same $\alpha^{n-1}$-slope. Then we can prove that, for $j=1,2$, $h_{j}=\oplus_{i=1}^{k}(h_{j}|_{\mathcal{E}_{i}})$ and each $h_{j}|_{\mathcal{E}_{i}}$ defines a $T$-adapted HYM metric on $\mathcal{E}_{i}$ (Theorem 3.29). Hence, by the uniqueness result for $\alpha^{n-1}$-slope stable sheaves, we can see that $h_{1}=h_{2}$.
@@ Acknowledgment
The author would like to thank his supervisor Ryushi Goto for his helpful comments and nice discussions. The author also thanks Prof. Yoshinori Hashimoto for suggesting the Uhlenbeck compactness to solve the problem and discussions. The author would like to thank Prof. Yuji Odaka for introducing him to the important paper [CCHTST25]. The author is grateful to Prof. Yuta Watanabe and Rei Murakami for fruitful discussions and important comments.
@@ 3. $(3)$
Let $X$ be a compact normal variety in Fujiki’s class $\mathcal{C}$ and $\alpha\in H^{1,1}_{BC}(X)$ be a pseudo-effective class on $X$. Let $\pi:Y\to X$ be a resolution of singularities of $X$ so that $Y$ is smooth Kähler. Then, we define the non-Kähler locus of $\alpha$ by
@@
Let $X$ be a compact normal analytic variety and $\alpha\in H^{1,1}_{BC}(X)$ be a nef and big class on $X$. Let $\pi:Y\to X$ be a resolution of singularities of $X$ so that $Y$ is a compact Kähler manifold. Then a torsion free sheaf $\mathcal{E}$ on $X$ is $\alpha^{n-1}$-slope stable (resp. semistable, polystable) if $\pi^{*}\mathcal{E}/{\rm{Tor}}$ is $(\pi^{*}\alpha)^{n-1}$-slope stable (resp. semistable, polystable).
@@ 2. $(3)$
We define adapted currents in singular settings via resolutions as follows.
Definition 2.18.
Let $X$ be a compact normal analytic variety and $\alpha\in H^{1,1}_{BC}(X)$ be a nef and big class. Let $\pi:Y\to X$ be a resolution of singularities so that $Y$ is a smooth Kähler manifold with a Kähler class $\omega_{Y}$. Then, a closed positive $(1,1)$-current $T$ in $\alpha$ is adapted if $\pi^{*}T$ is an adapted closed positive $(1,1)$-current in $\pi^{*}\alpha$ on $Y$ in the sense of Definition 2.15
The following lemma shows that the above definition 2.18 is independent of the choices of resolution $\pi$. The proof is obvious.
@@
Let $X$ be a compact normal Kähler variety with log terminal singularities. Assume that $X$ admits a singular Kähler-Einstein metric $\omega$. Then $\omega$ is an adapted current in the sense of Definition 2.18.
@@ 3. $(3)$
We can define $T$-adapted HYM metrics in singular settings via resolution as follows. The readers can find the definition of the ample locus of big classes in singular settings from Definition 2.3.
Definition 2.22.
Let $X$ be a compact normal analytic variety and $\alpha\in H^{1,1}_{BC}(X)$ be a nef and big class. Let $\mathcal{E}$ be a torsion free sheaf on $X$. Let $\pi:Y\to X$ be a resolution of singularities so that $Y$ is smooth Kähler and $E:=\pi^{*}\mathcal{E}/{\rm{Tor}}$ is locally free. Then, a $T$-adapted Hermitian-Yang-Mills metric on $\mathcal{E}$ is a smooth hermitian metric $h$ on the locally free locus of $\mathcal{E}|_{{\rm{Amp}}(\alpha)}$ such that $\pi^{*}h$ is a $\pi^{*}T$-adapted HYM metric on $E$.
By the following lemma, we can see that the Definition 2.22 is independent of the choice of a resolution $\pi$. The proof is obvious.
@@
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class. Let $T$ be a closed positive $(1,1)$-current in $\alpha$ which is smooth Kähler on ${\rm{Amp}}(\alpha)$. Let $E$ be a holomorphic vector bundle on $X$. Let $\pi:Y\to X$ be a composition of blow-ups along complex submanifolds contained in $E_{nK}(\alpha)$ or of codimension at least 2. Then a smooth hermitian metric $h$ on $E|_{{\rm{Amp}}(\alpha)}$ is a $T$-adapted HYM metric on $E$ if and only if $\pi^{*}h$ is a $\pi^{*}T$-adapted HYM metric on $\pi^{*}E$. In particular, the Definition 2.22 is independent of the choice of a resolution $\pi$.
@@ 3. $(4)$
  4. $(5)$
    There exists an analytic subset $\Sigma$ in $X\setminus D$ of codimension at least 2 and a $T$-admissible HYM connection $\nabla_{\infty}$ on $(E,h_{0})|_{X\setminus Z}$, where $Z=\overline{D\cup\Sigma}$, such that $A_{i}:=\nabla_{i}-\nabla_{h_{0}}$, a smooth ${\rm{End}}(E)$ valued 1-form, satisfies $\|A_{i}\|_{C^{k}(A,\omega_{i},h_{0})}\to 0$ in $i\to\infty$ for any $k>0$ and any compact subset $A\subset X\setminus Z$.
In the following subsections, we construct $\Psi_{\infty}^{\pm 1}$ the nontrivial $L^{2}_{1}(X\setminus Z,T)$-local limit of $\Psi_{i}^{\pm 1}$ in Lemma 3.5 above, which gives a $T$-adapted HYM metric on $E$ (refer to Theorem 3.17).

### 18 Mar 2026 v2

@@ 1 Introduction
The Kobayashi-Hitchin correspondence asserts that a holomorphic vector bundle $E$ over a compact complex manifold $X$ is slope polystable with respect to a Kähler class $\{\omega\}$ on $X$ if and only if $E$ admits an $\omega$-Hermitian-Yang-Mills metric. If $\omega$ is a smooth Kähler metric, this correspondence was proved by Donaldson [Don85] and Uhlenbeck-Yau [UY86] for holomorphic vector bundles on compact Kähler manifolds, by Bando-Siu [BS94] for reflexive sheaves on compact Kähler manifolds by introducing the notion of admissible Hermitian-Yang-Mills metrics, and by Xuemiao Chen [Chen25] for reflexive sheaves on compact normal Kähler varieties. More recently, the correspondence has been extended to compact klt Kähler varieties endowed with orbifold Kähler metrics motivated by the study of the equality cases of the Bogomolov-Gieseker inequality and the Miyaoka-Yau inequality of klt varieties [Faulk22], [CGNPPW23], [FO25] (refer to the introduction of [IJZ25] for more detail). There are also results when Kähler metrics have conical singularities [Li00], [JL25].
@@ 1.1.1 Definitions
Our main result of this paper is Theorem 1.4. To formulate Theorem 1.4, we firstly introduce the notion of adapted closed positive $(1,1)$-currents, inspired by the notion of tame approximations of [CCHTST25] for closed positive $(1,1)$-currents with bounded potentials.
Definition 1.1 (=Definition 2.15).
@@ 3. $(3)$
We will see that any nef and big class contains an adapted current (Proposition 2.16). One of the important examples of adapted currents is a singular Kähler-Einstein metric on a compact normal Kähler variety with log terminal singularities (Theorem 2.19). We next introduce the notion of adapted Hermitian-Yang-Mills metrics as admissible Hermitian-Yang-Mills metrics of Bando-Siu [BS94] with additional two conditions, called as adapted conditions:
Definition 1.2 (=Definition 2.20).
@@ 2. $(2)$
Let $X$ be a compact normal analytic variety which is bimeromorphic to a Kähler manifold and $\alpha\in H^{1,1}_{BC}(X)$ be a nef and big class. Let $\mathcal{E}$ be a reflexive sheaf on $X$. Then the following conditions are equivalent.
@@
    Let $\pi:Y\to X$ be any resolution of singularities of $X$ such that $Y$ is a compact Kähler manifold, $D=E_{nK}(\pi^{*}\alpha)$ is a snc divisor and $\pi^{*}\mathcal{E}/{\rm{Tor}}$ is a locally free sheaf. Let $T\in\pi^{*}\alpha$ be an adapted current. Then $\pi^{*}\mathcal{E}/{\rm{Tor}}$ admits a $T$-adapted Hermitian-Yang-Mills metric.
Furthermore, if $\pi^{*}\mathcal{E}/{\rm{Tor}}$ admits a $T$-adapted Hermitian-Yang-Mills metric, then it is unique up to constant multiplications.
@@
By [CCHTST25], it has proved that the pull-back of a singular Kähler-Einstein metric on a compact klt variety along a resolution of singularities is an adapted current. Hence we obtain the following result as a direct consequence of Corollary 1.6. Following Corollary 1.8 itself is a new result to the best of our knowledge.
@@
    Let $\pi:Y\to X$ be a resolution of singularities of $X$ such that $Y$ is a compact Kähler manifold, $D=E_{nK}(\pi^{*}\alpha)$ is a snc divisor and $\pi^{*}\mathcal{E}/{\rm{Tor}}$ is a locally free sheaf. Then $\pi^{*}\mathcal{E}/{\rm{Tor}}$ admits an $\pi^{*}\omega$-adapted Hermitian-Yang-Mills metric.
Furthermore, if $\pi^{*}\mathcal{E}/{\rm{Tor}}$ admits an $\pi^{*}\omega$-adapted Hermitian-Yang-Mills metric, then it is unique up to constant multiplications.
@@ 1.1.3 A priori estimates and applications
Corollary 1.9 (=Theorem 3.17, Corollary 3.28).
@@ 3. $(3)$
Let $X$ be a compact normal analytic variety which is bimeromorphic to a compact Kähler manifold and $\alpha\in H^{1,1}_{BC}(X)$ be a nef and big class. Then, for any $\alpha^{n-1}$-slope polystable reflexive sheaves $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$, their direct sum $\mathcal{E}_{1}\oplus\mathcal{E}_{2}$, tensor product $\mathcal{E}_{1}\otimes\mathcal{E}_{2}$ and wedge product $\mathcal{E}_{1}\wedge\mathcal{E}_{2}$ are $\alpha^{n-1}$-slope polystable.
@@ 1.1.4 Further discussions
If a nef and big class $\alpha$ is a Kähler class, then above map is bijective and it is the non-abelian Hodge correspondence in the case that a Higgs structure is trivial (refer to [ZZZ26] for more detailed and precise statement). Then it is a natural question whether the above map is bijective in this nef and big class-setting. The identification of the image of the above map will be important. To generalize the representation from $PU(n)$ above to $PGL(n+1,\mathbb{C})$, we will need to generalize the main Theorem 1.4 of this paper to holomorphic Higgs bundles.
@@
$\sup_{X}|s_{D}\Psi_{\infty}|_{h_{D}\otimes h_{0}\otimes h_{0}^{*}}\leq C$ (1.3)
where $h_{D}$ is a smooth hermitian metric on $\mathcal{O}(D)$ such that $|s_{D}|_{h_{D}}\leq 1$ (refer to Lemma 3.8). In fact, such $\Psi_{\infty}$ defines a hermitian metric $h_{0}\Psi_{\infty}^{2}$ on $E|_{X\setminus D}$ and it gives a $T$-adapted HYM metric (refer to Theorem 3.14 and Theorem 3.17). To establish the uniform $C^{0}$-estimate in (1.3), the mean value type inequality of Guo-Phong-Sturm [GPS24] (Lemma 3.6) plays an important role (refer to Remark 1.14). In fact, we can easily shows
@@
holds with universal constant $C>0$ (Lemma 3.9). Then, we can prove that if the $L^{1}$ norm in LHS in (1.4), denoted by $l_{i}$, diverges to $\infty$ in $i\to\infty$, then the normalized morphism $u_{i}:=l_{i}^{-1}s_{i}$ converges to a nontrivial limit $u_{\infty}$ with ${\rm{rk}}(u_{\infty})<{\rm{rk}}E$ (Lemma 3.10). Then, as the argument in [Sim88, Section 5], we can see that the projections to subspaces of the kernel of $u_{\infty}$ define $T$-weak holomorphic subbundles of $E$. Then, using the Chern-Weil formula for $T$-weak holomorphic subbundles (Proposition 1.5), we can see that some of the projections produce a destabilizing subsheaf of $E$ (Lemma 3.10). It contradicts to the $\alpha^{n-1}$-slope stability of $E$. Hence the $L^{1}$-norm of (1.4) is uniformly bounded and then $\Psi_{i}$ and $\Psi_{i}^{-1}$ converge to nontrivial morphisms $\Psi_{\infty}$ and $\Psi_{\infty}^{-1}$ respectively satisfying (1.3) (Lemma 3.8 and Lemma 3.9). Then we can see that $\Psi_{\infty}$ gives a $T$-adapted HYM metric $h_{0}\Psi_{\infty}^{2}$ on $E$ (refer to Theorem 3.14 and Theorem 3.17).
Remark 1.14.
To the best of the author’s knowledge, the mean value type inequality of Guo-Phong-Sturm [GPS24] was first applied in the context of Yang-Mills connections by J. Cao, P. Graf, P. Neumann, M. Păun, T. Peternell and X. Wu [CGNPPW23, section 3]. One of the major differences from their argument is that we apply the mean value type inequality of [GPS24] not to $\log|\Psi_{i}|^{2}=\log{\rm{Tr}}(\Psi_{i}^{2})$, but to $\log|s_{D}\Psi_{i}|^{2}$. This difference is essential in our argument. Indeed, since $T$, the limit of the sequence $\{\omega_{i}\}_{i}$, is not positive definite along $D$, the contraction $|\Lambda_{\omega_{i}}F_{h_{0}}|$ of the curvature of the initial hermitian metric $h_{0}$ is not uniformly bounded around $D$. Therefore, to apply the mean value type inequality of [GPS24], it is necessary that the closure of $\Omega_{a}=\{\log|s_{D}\Psi_{i}|^{2}+a>0\}$ be contained in a compact subset of $X\setminus D$. We also remark that the necessity of multiplying by $s_{D}$ is the main source of difficulties in this paper.
@@
for some $l\in\mathbb{Z}_{>0}$ where $h_{D}$ is a smooth hermitian metric on $\mathcal{O}(D)$ with $|s_{D}|_{h_{D}}\leq 1$. The technical lemma is to show that $g$ is holomorphic, that is, the following holds (Claim 3.25)
@@
on $X\setminus D$. [2nd step] Then, by Proposition 1.10, we obtain that a $T$-adapted HYM metric on an $\alpha^{n-1}$-slope stable vector bundle is unique (Lemma 3.24). [3rd step] Then, by applying the existence of $T$-adapted HYM metrics on $\alpha^{n-1}$-slope stable sheaves and 2nd step, we can prove that a holomorphic vector bundle $E$ admitting a $T$-adapted HYM metric is $\alpha^{n-1}$-slope polystable (Theorem 3.26). [4th step] Let $E|_{X\setminus D}=(\bigoplus_{i=1}^{k}\mathcal{E}_{i})|_{X\setminus D}$ be the decomposition of $E$ into $\alpha^{n-1}$-slope stable torsion free sheaves of same $\alpha^{n-1}$-slope. Then we can prove that, for $j=1,2$, $h_{j}=\oplus_{i=1}^{k}(h_{j}|_{\mathcal{E}_{i}})$ and each $h_{j}|_{\mathcal{E}_{i}}$ defines a $T$-adapted HYM metric on $\mathcal{E}_{i}$ (Theorem 3.27). Hence, by the uniqueness result for $\alpha^{n-1}$-slope stable sheaves, we can see that $h_{1}=h_{2}$.
@@ Acknowledgment
The author would like to thank his supervisor Ryushi Goto for his helpful comments and nice discussions. The author also thanks Prof. Yoshinori Hashimoto for essential comments and discussions. The author would like to thank Prof. Yuji Odaka for introducing him to the important paper [CCHTST25]. The author is grateful to Prof. Yuta Watanabe and Rei Murakami for fruitful discussions and important comments.
@@ 3. $(3)$
Let $X$ be a compact normal variety which is bimeromorphic to a compact Kähler manifold and $\alpha\in H^{1,1}_{BC}(X)$ be a pseudo-effective class on $X$. Let $\pi:Y\to X$ be a resolution of singularities of $X$ so that $Y$ is smooth Kähler. Then, we define the non-Kähler locus of $\alpha$ by
@@
Let $X$ be a compact normal analytic variety which is bimeromorphic to a compact Kähler manifold and $\alpha\in H^{1,1}_{BC}(X)$ be a nef and big class on $X$. Let $\pi:Y\to X$ be a resolution of singularities of $X$ so that $Y$ is a compact Kähler manifold. Then a torsion free sheaf $\mathcal{E}$ on $X$ is $\alpha^{n-1}$-slope stable (resp. semistable, polystable) if $\pi^{*}\mathcal{E}/{\rm{Tor}}$ is $(\pi^{*}\alpha)^{n-1}$-slope stable (resp. semistable, polystable).
@@ 2. $(3)$
The following lemma will be used later. The proof is obvious.
@@
Let $X$ be a compact normal Kähler variety with log terminal singularities. Assume that $X$ admits a singular Kähler-Einstein metric $\omega$. Let $\pi:Y\to X$ be a resolution of singularities such that $Y$ is a compact Kähler manifold. Then $\pi^{*}\omega$ is an adapted current in the sense of Definition 2.15.
@@ 3. $(3)$
The following lemma will be used later. The proof is obvious.
@@
Let $X$ be a compact Kähler manifold and $\alpha$ be a nef and big class. Let $T$ be a closed positive $(1,1)$-current in $\alpha$ which is smooth Kähler on ${\rm{Amp}}(\alpha)$. Let $E$ be a holomorphic vector bundle on $X$. Let $\pi:Y\to X$ be a composition of blow-ups along complex submanifolds contained in $E_{nK}(\alpha)$ or of codimension at least 2. Then a smooth hermitian metric $h$ on $E|_{{\rm{Amp}}(\alpha)}$ is a $T$-adapted HYM metric on $E$ if and only if $\pi^{*}h$ is a $\pi^{*}T$-adapted HYM metric on $\pi^{*}E$.
@@ 3. $(4)$
In the following subsections, we construct $\Psi_{\infty}^{\pm 1}$ the nontrivial limit of $\Psi_{i}^{\pm 1}$ in Lemma 3.5 above, which gives a $T$-adapted HYM metric on $E$ (refer to Lemma 3.7, Lemma 3.8, Lemma 3.9, Lemma 3.10 and Theorem 3.17).